# Navigating the Path of Writing: Outline-guided Text Generation with Large Language Models

### v1

@@ 6.2 Outline Evaluation
Automatic Evaluation To see the effects of the outline augmentation module, we conducted automatic evaluations on the initial and augmented outlines using criteria described in Section 4.1.1. The results in Table 2 show significant improvements in Coherence and Repetition aspects for the augmented outlines compared to the initial ones, indicating that the outline augmentation process enhances content consistency and reduces unnecessary repetition. Notably, although Diversity and Coherence are often considered trade-offs, the augmented outlines in our study maintained Diversity while improving Coherence. This suggests that the outline expansion module can increase consistency without compromising content diversity. Detailed performance across various text domains is in Appendix 3.
Human Evaluation As described in Section 4.1.2, we conducted human evaluations to assess the cohesion, natural flow, diversity, and redundancy of initial and augmented outlines. The augmented outlines demonstrated significant improvements in all aspects except cohesion, which slightly declined or remained stable. Nevertheless, the overall performance of the augmented outlines surpassed that of the initial outlines. Further evaluations of the augmented outlines were conducted on usefulness and improvement, which indicated the extent of useful information added and overall quality enhancement compared to the initial outlines. As shown in Figure 8, all models demonstrated improvements in both metrics, validating the power of the browsing step. Detailed performance across various text domains is in Appendix 4.
@@ 7 Conclusion
We introduced Writing Path, a novel framework that enhances the ability of LLMs to generate high-quality and goal-oriented writing by employing explicit outlines. Our approach emphasizes the importance of providing structured guidelines to control the quality of the generated text from the early stages of the writing process. The Writing Path framework consists of five stages: prepare metadata, generate title and initial outline, browse for information, generate augmented outline, then write the text.
We verified the impact of Writing Path by conducting a comprehensive evaluation that incorporates automatic and human evaluations covering a wide range of aspects. Our experimental results demonstrate that texts generated following the full Writing Path approach, which includes the use of augmented outlines, exhibit superior performance compared to texts produced using only initial outlines or without any intermediate outlines. We also proposed a framework for assessing the Writing Path’s intermediate outlines, which found that augmented outlines have better inherent quality than initial outlines, demonstrating the importance of outline augmentation steps.
@@ 8 Acknowledgments
Appendix A Prompts
Writing Evaluation Prompt
@@ Appendix A Prompts
You will be given one text written for a blog post.
@@
Keep this document open while reviewing and refer to it as needed.
You are a writing expert! it is crucial to apply a robust evaluation.
Evaluation Criteria:
@@
Table 8: Writing Evaluation Prompt for Checklist-based Assessment.
Prompt for Metadata construction (step #1)
@@
1. the purpose of the post, 2. the type of post, 3. the style of the post, and 4. keywords.
@@
Prompt for Generation of Title and Initial Outline(step #2)
@@
Following the example above, based on the post information provided below,
only create "==Title==" and a brief "==Initial Outline==".
@@
Prompt for Generation of Augmented Outline (step #4)
Map the necessary additional information below to create an augmented outline. Here is an example.
@@
Prompt for Generation of Text (step #5)
@@
Table 9: Writig Path Prompt for Each Stage.

### v2

@@ 5.2 Outline Evaluation
Automatic Evaluation To see the effects of the outline augmentation module, we conducted automatic evaluations on the initial and augmented outlines using criteria described in Section 3.1.1. The results in Table 1 show significant improvements in Coherence and Repetition aspects for the augmented outlines compared to the initial ones, indicating that the outline augmentation process enhances content consistency and reduces unnecessary repetition. Notably, although Diversity and Coherence are often considered trade-offs, the augmented outlines in our study maintained Diversity while improving Coherence. This suggests that the outline expansion module can increase consistency without compromising content diversity. Detailed performance across various domains is in Table 3.
Human Evaluation As described in Section 3.1.2, we conducted human evaluations to assess the cohesion, natural flow, diversity, and redundancy of initial and augmented outlines. The augmented outlines demonstrated significant improvements in all aspects except cohesion, which slightly declined or remained stable. Nevertheless, the overall performance of the augmented outlines surpassed that of the initial outlines. Further evaluations of the augmented outlines were conducted on usefulness and improvement, which indicated the extent of useful information added and overall quality enhancement compared to the initial outlines. As shown in Figure 6, all models demonstrated improvements in both metrics, validating the power of the browsing step. Detailed performance across various text domains is in Table 4.
6 Real-World Deployment
WritingPath was integrated into a commercial blogging platform as a writing assistance feature and tested for six months. In the service environment, additional considerations such as safety filtering and content quality control measures were necessary for reliable content generation. The system architecture of CLOVA for Writing by NAVER is depicted in Figure 7.
Figure 7: Real-world deployment pipeline of WritingPath.
The serving pipeline integrates multiple components for reliable service operation. It integrates user request handling, content filtering, Kafka pipeline, and retrieval. Requests pass through a Gateway with rate limiting and are filtered for harmlessness. Specifically, the system includes emergency filtering and safety classification before passing requests to WritingPath. Additionally, a token event monitoring system tracks model usage, followed by token event filtering over output anomalies.
@@ 7 Conclusion
We introduced WritingPath, a framework that enhances the ability of LLMs to generate high-quality and goal-oriented writing by employing explicit outlines. Designed for real-world content creation, our approach uses structured guidelines from the early stages to ensure consistent quality control.
We verified the impact of WritingPath by conducting a comprehensive evaluation that incorporates automatic and human evaluations covering a wide range of aspects. Our experimental results demonstrate that texts generated following the full WritingPath approach, which includes the use of augmented outlines, exhibit superior performance compared to texts produced using only initial outlines or without any intermediate outlines. We also proposed a framework for assessing the WritingPath’s intermediate outlines, which found that augmented outlines have better inherent quality than initial outlines, demonstrating the importance of outline augmentation steps. We hope that this work will contribute to the research and development of more reliable AI-assisted writing solutions.
@@ 8 Acknowledgments
Appendix A Related Work
A.1 Collaborative Writing with Language Models
Recursive Reprompting and Revision technique extends the potential of LMs to assist not only with draft generation but also with editing and revision processes. This approach employs LLM prompt chains such as planning - drafting - reviewing - suggesting revisions in an iterative fashion to enhance the quality of written content. Re3 (Yang et al., 2022) introduces a framework for maintaining the long-range coherence of draft generation. It operates separate rewriter and edit modules in its prompt chain to check and refine plot relevance and long-term factual consistency. PEER (Schick et al., 2023) proposes a recursive revision framework based on the concept of self-training, where the model autonomously selects the editing operations for revision and provides explanations for the modifications it makes. RECURRENTGPT Zhou et al. (2023) utilizes a recursive, language-based mechanism to simulate LSTM Hochreiter and Schmidhuber (1997), enabling the generation of coherent and extended texts. While these works are relevant to collaborative writing with LMs, direct comparisons with our approach are unfeasible. These studies focus on specific tasks like story generation, requiring task-specific training and datasets, which are unavailable in Korean for our writing tasks.
A.2 Integrating External Information
A.3 Writing Evaluation
Appendix B Details of Dataset
We selected 20 blog posts for each domain [^7] https://blog.naver.com/, resulting in 100 seed data points. For seed data construction, we generated metadata, including purpose, topic, keywords, and expected reader, based on the title and content of the blog posts. This metadata is the input to the WritingPath, helping the model understand the context of the post and generate relevant outlines and text. We created a test dataset of 1,100 instances per model under evaluation using the seed data. Each data point includes the outputs of each WritingPath step: an outline, additional information, an augmented outline, and the final text. With analysis experiments as well, we generated a total of 1,500 posts for each model, resulting in 4,500 instances in total. For human evaluation, we randomly sampled 10% of the outlines and texts and assessed their scores. The final texts were evaluated by human experts, and the dataset aligns the generated outputs from three models with the human scores.
Appendix C Details of Evaluation
C.1 Compensation Details
Outline evaluators were compensated with a 6,000 KRW ($\approx$ 4.2 USD) gift card for their 30-minute participation. And writing experts were compensated at a rate of 9,000 KRW ($\approx$ 6.6 USD) per one-writing sample.
C.2 Automatic Evaluation - Outline
  • Logical alignment: Based on Chen and Eger (2023), we utilize Natural Language Inference (NLI) which examines whether the headers and subheaders within an outline logically connect, ensuring the structural integrity necessary for coherent argumentation888we utilize gpt-4-turbo for NLI evaluation.
binary binary binary binary binary binary binary binary
GPT-3.5 51.66 31.14 46.29 88.11 66.43 21.14 35.14 48.56
GPT-4 68.00 60.57 72.86 89.26 80.29 54.14 66.29 70.20
HyperCLOVA X 89.71 84.46 91.14 98.06 92.57 74.00 80.00 87.13
C.3 Human Evaluation - Outline
The human evaluation criteria are based on aspects considered in previous studies on text coherence, relevance, and quality assessment Yang et al. (2022, 2023); Zhou et al. (2023); Ke et al. (2022). For both initial and augmented outlines, the human evaluation is performed on the following five aspects, using a 1-4 point scale:
  • Natural Flow: Assesses whether the outline flows in a natural order.
Appendix D Further Analysis of Writing Quality
To further analyze the quality of the text generated through the complete WritingPath pipeline, we conducted a human evaluation based on the CheckEval framework. The results are presented in Table 2. The analysis by six writing experts showed that GPT-4 and HyperCLOVA X generally performed better than GPT-3.5 in terms of writing quality. HyperCLOVA X exhibited higher scores in specificity compared to other models, which is consistent with the findings reported in KMMLU Son et al. (2024) regarding the advantages of language-specific models. Detailed performance metrics across various domains and further LLM evaluations can be found in Table 56. Furthermore, We consider seven key aspects (Section 3) for evaluating the quality of writing. CheckEval’s binary responses for each aspect allow for identifying the specific factors contributing to the assessments. We found that logical fluency, coherence, consistency, and specificity significantly contribute to the improvement of text quality through the WritingPath (Figure 8).
During the evaluation of the writing quality, writing experts assigned binary overall quality ratings (1 for high quality, 0 for low quality) to the texts. We employed the Kendall tau correlation to examine the relationship between the overall binary ratings and the scores for each evaluation aspect. The analysis (Figure 9) revealed a significant correlation for all the aspects we designed. Interestingly, logical fluency, specificity, and coherence, which were found to be particularly important in determining the perceived quality of written content, are among the aspects that showed the most significant improvement through the WritingPath (Figure 8).
The progressive improvement in these aspects can be attributed to the effectiveness of using outlines. The initial outline (init) helps organize information more logically and coherently compared to using only metadata (meta), while the augmented outline (aug) further enhances the consistency and richness of the content. These findings highlight the importance of using outlines in the writing process and demonstrate how their gradual enhancement leads to better-structured, more coherent, and content-rich texts, ultimately improving the overall quality of the written output.
Figure 10: LLM Evaluation
Model Category Outline Type Logical Alignment Coherence Diversity Repetition
Metrics NLI ($\uparrow$) UCI ($\uparrow$) / NPMI ($\uparrow$) Topic Diversity ($\uparrow$) Self-BLEU ($\downarrow$)
Eval Level Header-Subheader Outline Outline Outline
GPT 3.5 Beauty Initial - 0.638 / 0.298 0.488 48.01
Augmented 0.483 1.506 / 0.553 0.513 23.79
Travel Initial - 0.835 / 0.454 0.708 21.56
Augmented 0.575 1.646 / 0.540 0.670 13.21
Gardening Initial - 0.496 / 0.206 0.591 26.52
Augmented 0.658 1.291 / 0.575 0.592 19.24
Cooking Initial - 0.543 / 0.352 0.641 15.94
Augmented 0.686 1.003 / 0.411 0.712 13.71
IT Initial - 0.491 / 0.235 0.523 25.67
Augmented 0.667 1.180 / 0.463 0.560 16.69
GPT 4 Beauty Initial - 0.908 / 0.574 0.657 36.50
Augmented 0.577 1.854 / 0.573 0.658 18.80
Travel Initial - 0.717 / 0.534 0.691 17.61
Augmented 0.615 1.690 / 0.530 0.688 10.63
Gardening Initial - 0.833 / 0.398 0.676 20.33
Augmented 0.724 1.559 / 0.555 0.681 13.43
Cooking Initial - 0.693 / 0.468 0.720 13.85
Augmented 0.701 1.512 / 0.464 0.745 10.98
IT Initial - 0.854 / 0.454 0.625 16.80
Augmented 0.702 1.448 / 0.471 0.633 11.77
HyperCLOVA X Beauty Initial - 1.030 / 0.629 0.810 22.37
Augmented 0.504 1.979 / 0.553 0.793 12.33
Travel Initial - 0.981 / 0.594 0.801 11.03
Augmented 0.626 2.285 / 0.590 0.843 10.20
Gardening Initial - 0.694 / 0.280 0.623 20.73
Augmented 0.693 1.833 / 0.563 0.624 12.87
Cooking Initial - 0.526 / 0.251 0.603 17.13
Augmented 0.774 1.416 / 0.454 0.658 8.99
IT Initial - 0.528 / 0.277 0.606 19.22
Augmented 0.776 1.560 / 0.536 0.596 13.13
Table 3: Detailed outline automatic evaluation results.
Model Category Outline Type Cohesion Natural Flow Diversity Redundancy Usefulness Improvement
GPT 3.5 Beauty Initial 3.542 2.958 2.833 2.917 - -
Augmented 3.208 2.875 3.417 3.375 3.000 0.542
Travel Initial 3.625 2.833 3.167 2.917 - -
Augmented 3.708 3.125 3.750 3.542 3.458 0.708
Gardening Initial 2.958 2.375 2.542 2.417 - -
Augmented 3.042 2.833 3.375 2.667 2.542 0.708
Cooking Initial 3.292 2.417 2.417 2.583 - -
Augmented 2.708 2.333 3.458 2.833 2.542 0.458
IT Initial 3.458 2.917 2.875 2.792 - -
Augmented 3.083 2.750 3.708 3.250 2.875 0.542
GPT 4 Beauty Initial 3.375 2.750 3.083 3.167 - -
Augmented 3.542 3.208 3.708 3.583 3.208 0.833
Travel Initial 3.542 2.792 3.042 2.833 - -
Augmented 3.625 3.083 3.792 3.542 3.333 0.750
Gardening Initial 3.792 3.125 3.083 2.917 - -
Augmented 3.625 3.208 3.833 3.500 3.208 0.667
Cooking Initial 3.292 2.708 2.833 2.333 - -
Augmented 3.208 2.625 3.542 2.917 2.833 0.542
IT Initial 3.000 2.917 3.250 3.042 - -
Augmented 3.000 2.792 3.833 3.583 3.042 0.750
HyperCLOVA X Beauty Initial 3.375 3.292 2.583 3.125 - -
Augmented 3.500 3.667 3.958 3.833 3.667 0.917
Travel Initial 3.667 2.792 3.125 3.417 - -
Augmented 3.583 3.417 4.042 4.000 3.542 0.833
Gardening Initial 3.500 3.125 2.833 3.042 - -
Augmented 3.708 3.750 3.958 3.625 3.583 0.875
Cooking Initial 3.500 2.958 2.750 3.250 - -
Augmented 3.208 3.375 3.792 3.792 3.250 0.750
IT Initial 3.292 2.625 2.833 3.250 - -
Augmented 3.042 3.208 3.917 3.708 3.083 0.750
Table 4: Detailed outline human evaluation results.
Model Category Linguistic Fl. Logical Fl. Coh. Cons. Comple. Spec. Int. Overall
GPT 3.5 Beauty 96.00 75.42 90.00 94.17 89.58 52.50 63.06 80.10
Travel 100.00 78.54 97.08 96.94 97.08 65.83 81.67 88.16
Gardening 98.67 75.63 95.00 96.11 89.17 49.17 78.89 83.23
Cooking 99.50 72.92 95.42 97.78 95.00 35.83 84.17 82.94
IT 100.00 76.67 97.50 96.94 94.17 39.17 50.83 79.33
Total 98.83 75.83 95.00 96.39 93.00 48.50 71.72 82.75
GPT 4 Beauty 99.17 89.79 97.50 99.44 99.17 84.58 98.61 95.47
Travel 99.00 90.21 91.67 97.22 96.67 70.00 96.94 91.67
Gardening 99.67 90.00 94.17 98.33 100.00 74.17 97.78 93.44
Cooking 99.67 89.58 93.75 98.61 97.92 63.33 96.67 91.36
IT 99.17 88.96 88.33 96.11 99.17 41.67 76.11 84.22
Total 99.33 89.71 93.08 97.94 98.58 66.75 93.22 91.23
HyperCLOVA X Beauty 100.00 88.33 98.33 100.00 90.00 90.42 91.38 94.07
Travel 99.50 83.75 90.42 97.50 91.25 79.58 92.77 90.68
Gardening 99.33 88.13 93.33 98.61 95.00 70.42 84.16 89.85
Cooking 98.67 82.29 88.75 96.67 90.42 87.92 91.11 90.83
IT 98.50 76.04 81.25 92.50 87.08 55.00 52.50 77.55
Total 99.20 83.71 90.42 97.06 90.75 76.67 82.38 88.60
Table 5: Detailed writing LLM evaluation results.
Model Category Linguistic Fl. Logical Fl. Coh. Cons. Comple. Spec. Int. Overall
GPT 3.5 Beauty 51.43 30.29 51.43 92.57 67.14 14.29 29.29 48.06
Travel 68.00 56.00 68.57 87.43 82.14 47.14 55.71 66.43
Gardening 45.14 30.86 44.29 87.43 52.86 18.57 30.00 44.16
Cooking 46.29 8.86 21.43 80.00 72.86 7.14 32.86 38.49
IT 47.43 29.71 45.71 93.14 57.14 18.57 27.86 45.65
Total 51.66 31.14 46.29 88.11 66.43 21.14 35.14 48.56
GPT 4 Beauty 72.00 67.43 82.86 93.71 86.43 56.43 72.14 75.86
Travel 76.00 71.43 82.86 89.71 86.43 67.14 79.29 78.98
Gardening 66.29 61.14 75.71 85.71 70.71 54.29 61.43 67.90
Cooking 63.43 50.86 61.43 90.86 82.14 47.14 62.14 65.43
IT 62.29 52.00 61.43 86.29 75.71 45.71 56.43 62.84
Total 68.00 60.57 72.86 89.26 80.29 54.14 66.29 70.20
HyperCLOVA X Beauty 92.00 87.43 91.43 99.43 92.14 69.29 81.43 87.59
Travel 95.43 91.43 95.71 99.43 100.00 84.29 85.00 93.04
Gardening 88.57 85.71 97.14 99.43 95.71 75.71 82.86 89.31
Cooking 90.86 85.71 90.00 98.29 96.43 81.43 82.14 89.27
IT 81.71 72.00 81.43 93.71 78.57 59.29 68.57 76.47
Total 89.71 84.46 91.14 98.06 92.57 74.00 80.00 87.13
Table 6: Detailed writing human evaluation results.
Aspect Subaspect Descriptions
Natural Expression
Does the given text read naturally without any unnatural rhythm or excessively emphasized parts?
주어진 글이 부자연스러운 리듬이나 과도하게 강조된 부분 없이 자연스럽게 읽히나요?
Text Length
Is the length of the text suitable for the purpose and is it not excessively verbose or overly concise?
텍스트의 길이가 목적에 적합하며 과도하게 장황하거나 지나치게 간결하지는 않은 글인가요?
Vocabulary
Is the vocabulary appropriate for the context, not overly complex, and suitable for the topic and reader?
어휘가 맥락에 맞지 않거나 지나치게 복잡하지 않고, 주제와 독자에 적합한가요?
Syntax
Is the composition and sentence structure of the given text correct?
주어진 글의 구성과 문장의 구조가 올바른가요?
Linguistic Fluency Mechanic-Spelling, Punctuation
Is the spelling and punctuation of the given text correctly applied?
주어진 글의 철자와 문장부호가 올바르게 적용되었나요?
Organization (layout)
Does the given text have a clear and effective structure (layout)?
주어진 글은 명확하고 효과적인 구조 (레이아웃)를 가지고 있나요?
Repetitive Content
Is the text free of repetitive or unnecessary content?
텍스트 내에서 반복되는 내용이나 불필요한 내용이 없는 글인가요?
In the text, are the sentences well connected and progressing naturally and logically?
글 내에서 문장들이 잘 연결되어 있어 자연스럽고 논리적으로 진행이 되나요 ?
Inter-sentence Cohesion
Did you use conjunctions appropriately to improve readability?
가독성을 높이기 위한 접속사를 적절하게 사용했나요?
Logical Fluency Inter-paragraph Cohesion
Are the paragraphs in the text logically connected and progressing with each other?
텍스트 내의 단락들이 논리적으로 연결되어 서로 진행되나요?
Topic Consistency
Is the entire article consistently progressing with the central theme as the focus?
전체 글이 중심 주제를 중심으로 일관되게 진행되나요?
Coherence Topic Setence and Paragraph
Does each paragraph of the article have a clear subtopic centered around the main idea?
글의 각 문단이 주요 아이디어를 중심으로 명확한 소주제를 가지고 있나요?
Is a consistent narrative tone and style maintained throughout the entire text?
텍스트 전체에서 일관된 서술 어조와 어투가 유지되나요?
Tone
Is there no sudden change in tone in the context of the writing?
글의 맥락에서 급격한 어조 변화가 없는 글인가요?
Stance/Posture
Does the author present a consistent opinion on the topic in the writing? (Should not present conflicting opinions on the same subject)
저자는 글에서 주제에 대한 일관된 의견을 제시하나요? (동일한 대상에 대한 상반된 의견을 제시하지 않아야 함)
Does the given text maintain a consistent style type (spoken language, written language, informal, formal, etc.)?
주어진 글이 일관된 스타일의 유형 (구어체, 문어체, 반말, 존댓말 등의 유형)을 유지하나요?
Consistency Style
Do you consistently use abbreviations and acronyms when necessary?
필요 시 약어와 머리글자가 일관되게 사용되나요?
Is it a clear text that does not excessively use uncommon or complex words?
일반적이지 않거나 복잡한 단어들이 과도하게 등장하지 않는 명료한 글인가요?
Vocabulary
Is the definition of unfamiliar and difficult words provided and are they used appropriately in context?
낯설고 어려운 단어에 대한 정의가 되어 있고 문맥에 맞게 잘 사용되었나요?
Is the given text clearly structured without excessively complex sentence structures?
주어진 글이 과도하게 복잡한 문장 구조를 가진 문장들 없이 명확하게 구성되어 있나요?
Complexity Syntax
Do the first sentences of each paragraph start differently? (Asking if the text has paragraphs that do not all start the same way)
각 문단의 첫 문장이 다양하게 시작되나요? (각 문단의 시작이 모두 동일하지 않은 글인지 질문)
Is the example appropriately connected to the topic of the article?
예시가 글의 주제와 적절하게 연결되어 있나요?
Use of Examples and Review
Was the author’s personal experience mentioned specifically?
작성자의 개인적인 경험이 구체적으로 언급되었나요?
In the writing, were specific numerical values such as ratios and quantities mentioned?
글에서 구체적으로 비율, 수량과 같은 수치들이 언급되었나요?
Specificity Detailed Descriptions
When introducing details in a writing, do you appropriately utilize context or background information?
글에서 세부 사항을 소개할 때 맥락이나 배경 정보를 적절하게 활용하나요?
Engagement
Was the blog post written based on an appealing storytelling approach? (It’s okay if an exaggerated tone is included)
블로그 글이 매력적인 스토리텔링 접근 방식을 기반으로 작성되었나요? (과장된 어조가 포함되어도 괜찮음)
Kindness
Was the written blog post written in a friendly tone for the readers?
작성된 블로그 글은 독자들에게 친근한 어조로 작성되었나요?
Does the written blog post include the author’s unique ideas or perspectives?
작성된 블로그 글에는 작성자의 독특한 아이디어나 관점이 포함되어 있나요?
Interestingness Originality
Does the writer’s personal experience add freshness to the writing?
작성자의 개인적인 경험이 글에 신선함을 더하나요?
Table 7: Evaluation principles.
@@ Appendix D Further Analysis of Writing Quality
Writing Evaluation Prompt You will be given one text written for a blog post.
@@
Keep this document open while reviewing and refer to it as needed. You are a writing expert! it is crucial to apply a robust evaluation.
## Evaluation Criteria:
@@
Figure 11: Writing Evaluation Prompt for Checklist-based Assessment.
WritingPath Prompt Prompt for Metadata construction (step #1):
@@
1. the purpose of the post
2. the type of post
3. the style of the post
4. keywords.
@@
Prompt for Generation of Title and Initial Outline (step #2):
@@
Following the example above, based on the post information provided below, only create "==Title==" and a brief "==Initial Outline==".
@@
Figure 12: WritingPath Prompt for Each Stage (Step 1 and 2).
WritingPath Prompt Prompt for Generation of Augmented Outline (step #4): Map the necessary additional information below to create an augmented outline. Here is an example.
@@
Prompt for Generation of Text (step #5):
@@
Figure 13: WritingPath Prompt for Each Stage (Step 4 and 5).